OpenStudy (dragontears1128):

The football team has 30 more members than the basketball teams. If each group had 10 more members, the ratio of their membership would be 3:2. How many members are in each group? Write your answer as football x basketball y

OpenStudy (dragontears1128):

@AliSoliman

OpenStudy (alisoliman):

Football: 80 members basketball 50 members Assuming that 30+x for footbal team x for basketball team when adding 10 members more to each of them it becomes 40+x for football team x+10 for basketball team when the ratio between them Football : Basketball members = 3:2  then the equation is : 30 + 3x = 80 + 2x x= 50